OpenStudy (anonymous):

Let f(x) = -2x - 7 and g(x) = -4x + 3. Find (f ° g)(-5). A. 23 B. –53 C. –9 D. 3

hartnn (hartnn):

To find (fog)(x),u replace x in f(x) with entire g(x) expression,then replace x in the answer with -5 to get (fog)(-5)....try it

OpenStudy (anonymous):

I have no idea what your talking about. I cant do that. ha

hartnn (hartnn):

see,here i am replacing x in f(x) with g(x) f(g(x))=-2(-4x+3)-7=.....? now can u simplify this?

OpenStudy (anonymous):

-6-4x?

hartnn (hartnn):

f(g(x))=-2(-4x+3)-7=(-2)(-4x)+(-2)3-7=8x-6-7=8x-13.....now put x=-5

OpenStudy (anonymous):

Nevermind. I cant do all that!

hartnn (hartnn):

u have to just put x=-5 in 8x-13 like this 8(-5)-13= -40-13=-53.....there u go...

OpenStudy (anonymous):

Oh. I thought you were trying to give me some huge equation at first. thank you